OpenStudy (anonymous):

what is the proper way to set up this problem,a/b+c if a/b=5,a/c=2. I say the answer is 7?

OpenStudy (hoblos):

a/b = 5 so b/a = 1/5 a/c = 2 so c/a = 1/2 b/a + c/a = (b+c)/a = 1/2 + 1/5 = (5+2)/10 = 7/10 thus a/(b+c) = 10/7

OpenStudy (anonymous):

Sorry, I wrote it down wrong its a/b+c if a/b=5;b/c=2